Exactly 18.0 mL of water at 24.0 °C is added to a hot iron skillet. All of the water is converted to steam at 100.0°C. The mass of the skillet is 1.05 kg. What is the change in temperature of the skillet?
Answer #1

Mass of water = VOlume * density = 18 * 1.0 = 18 g

Heat added to water = mass of water * specific heat of water * Temperature change + mass of water * latent heat of vaporisation

= 18 * 4.184 * (100 - 24) + 18 *2260 = 46403.712 J

We have: heat released by iron = heat absorbed water = mass of skillet * Specific heat of iron * Temp. change

46403.712 = 1050 * 0.44 * Temp change

SOlving, temperature change = 100.44 degrees Celsius